Motor Homes Inc. (MHI) is presently enjoying abnormally high growth because of a surge in the demand for motor homes. The company expects free cash flow to grow at a rate of 20% for the next 4 years, after which there will be no growth (g = 0) in FCFF. The company’s last FCFF, FCFF0, was $1.50. The firm is consisted of 100% equity and has no debt. MHI’s beta is 1.5, the market risk premium is 6%, and the risk-free rate is 4%. Given there’s no debt, what is the current equity value of the firm using DCF model?
Answer #1

Here the valuation of equity is based on the concept of H-model , detailed as

Do(1+ gn) Do* H * (ga- gn) Ро (К - 9n) Po (К - 9n) _

P0 = Intrinsic value of share

D0= FCFF of Current year = 1.50

ga = Growth rate of FCFF at aggressive =20%

gn= Growth rate of FCFF at normalize rate =0%

H = n/2 = 4/2 = 2 years

K= Expected rate of return , (calculated by CAPM , K = risk free+Beta*Market risk premium)

K = 4%+ 1.5*6%=13%

P01.5 1.5 2 0.20 0.13 0.13 = $ 16.154